# Ergodic Risk Measures: Towards a Risk-Aware Foundation for Continual Reinforcement Learning

### v1

@@ 4.2 Ergodic Risk Measures for Continual Reinforcement Learning
Although a risk measure, $\rho$, is typically thought of as a functional that capture distributional characteristics of a random variable beyond its mean, if one were to set $\rho(X)=\mathbb{E}[X]$, then the notions of plasticity and local time consistency introduced in this work could be used as a formalism for the stability-plasticity dilemma in the risk-neutral continual setting.
@@ 5 Case Study: CVaR as an Ergodic Risk Measure
In this section, we present a case study in which we optimize an ergodic risk measure in a continual learning setting. In particular, we focus on optimizing the well-known conditional value-at-risk (CVaR) risk measure [Rockafellar and Uryasev, 2000]. More formally, consider a random variable $X$ with a finite mean on a probability space $(\Omega,\mathcal{F},\mathbb{P})$, and with a cumulative distribution function $F(x)=\mathbb{P}(X\leq x)$. The (left-tail) value-at-risk (VaR) of $X$ with parameter $\tau\in(0,1)$ represents the $\tau$-quantile of $X$, such that $\text{VaR}_{\tau}(X)=\sup\{x\mid F(x)\leq\tau\}$. When $F(x)$ is continuous at $x=\text{VaR}_{\tau}(X)$, $\text{CVaR}_{\tau}(X)$ can be interpreted as the expected value of $X$ conditioned on $X$ being less than or equal to $\text{VaR}_{\tau}(X)$, such that $\text{CVaR}_{\tau}(X)=\mathbb{E}[X\mid X\leq\text{VaR}_{\tau}(X)]$.
As per the results in Section 4, and given Assumptions 4.7 and 4.8, the CVaR risk measure can be formulated as the following continual learning objective:
$\text{CVaR}_{\pi_{t}}\doteq\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=t_{0}}^{n}\text{CVaR}[R_{t}\mid S_{t}\sim\mu_{\pi_{t}},A_{t}\sim\pi_{t}].$ (7)
That is, we want to optimize the (left-tail) conditional value-at-risk associated with the limiting per-step reward distribution induced when following stationary policy $\pi_{t}$. We note that, in addition to the CVaR objective (7) corresponding to an ergodic risk measure (as per the results of Section 4), it also corresponds to a coherent risk measure [Rockafellar and Uryasev, 2000].
In terms of the case study being presented, our aim is to optimize the CVaR objective (7) via the RED CVaR Q-learning algorithm proposed in Rojas and Lee [2025] in two continual learning tasks:
@@
In terms of empirical results, Figures 1 and 2 show the resulting agent behaviour as learning progresses in both tasks. In particular, Figure 1 shows that in the $\tau$-RPBP task, the agent correctly learns to stay in the blue world state in the beginning, and then correctly changes its preference to the red world state once its risk attitude changes from risk-neutral to risk-averse. Similarly, Figure 2 shows that in the $\mathcal{S}$-RPBP task, the agent is able to continually adapt and find the state with the better CVaR. The full set of experimental details and results can be found in Appendix B.
Figure 1: Rolling percent of time that the agent stays in the blue world state as learning progresses in the $\tau$-RPBP task. A solid line denotes the mean percent of time spent in the blue world state, and the corresponding shaded region denotes a 95% confidence interval over 50 runs. As shown in the figure, the agent correctly learns to stay in the blue world state in the beginning, and then correctly changes its preference to the red world state once its risk attitude changes from risk-neutral to risk-averse.
Figure 2: Rolling reward CVaR as learning progresses in the $\mathcal{S}$-RPBP task. A solid line denotes the mean CVaR, and the corresponding shaded region denotes a 95% confidence interval over 10 runs. The blue and red dashed lines denote the reward CVaR of the blue and red world states, respectively. As shown in the figure, the agent is able to continually adapt and find the state with the better CVaR.
@@ 6 Discussion
In this work, we took the first steps towards developing a risk-aware foundation for continual RL. In particular, we first examined the classical theory of risk measures, and showed that, in its current form, it is incompatible with continual RL, and particularly, the stability-plasticity dilemma. Then, building on this insight, we extended risk measure theory into the continual setting by introducing a new class of ergodic risk measures, which are designed to be compatible with continual learning. Finally, we provided a CVaR-based case study, along with numerical results, which showed the intuitive appeal and theoretical soundness of ergodic risk measures in a continual learning setting.
More broadly, the introduction of ergodic risk measures offers several potential benefits for the RL community. In particular, the introduction of the mathematical plasticity and local time consistency properties, which are at the heart of ergodic risk measures, effectively formalizes the stability-plasticity dilemma from the perspective of the optimization objective. Importantly, if one considers the risk-neutral case as a specific instance of the risk-aware case, then this formalization of the stability-plasticity dilemma could be applied more broadly in other continual RL settings. Moreover, in comparison to the static and nested (dynamic) risk measures that are used in the non-continual RL setting, we note that ergodic risk measures offer several advantages. In particular, ergodic risk measures retain some notion of time consistency, while remaining highly interpretable, thereby capturing the appeal of both static and nested risk measures.
@@ Appendix A Risk Measures
In essence, a dynamic risk measure provides a time-indexed family of risk assessments, allowing risk to be tracked and updated as new information becomes available. In RL-based contexts, dynamic risk measures can be useful for capturing the sequential nature of decision-making (i.e., that actions taken at each time step can potentially influence future outcomes, and hence, future risk evaluations).
@@
$\rho_{t+1}(X)\leq\rho_{t+1}(X^{\prime})\implies\rho_{t}(X)\leq\rho_{t}(X^{\prime}).$
Time-consistent risk measures can be appealing because they ensure that if one future outcome is deemed less risky than another at some time step, $t$, then that same outcome is not deemed more risky than the other at any other time step. In RL-based contexts, time-consistent risk measures can be useful because they can induce Bellman-like recursions with appealing dynamic programming-like properties. One way to think about time consistency in RL-based contexts is to ask the question: can the agent change its mind about how risky something is based on new information? If the answer is yes, then there exists a lack of time consistency.
@@
(Nested Risk Measure; adapted from Ruszczyński [2010]) A nested risk measure is a dynamic risk measure that is constructed recursively from one-step conditional risk measures. More formally, given conditional risk measures, $\rho_{t}:L^{\infty}(\mathcal{F}_{t+1})\to L^{\infty}(\mathcal{F}_{t})$, a nested risk measure over time horizon $N$ is defined as:
$\rho_{0:N}(X)\doteq\rho_{0}\big(\rho_{1}(\cdots\rho_{N-1}(X)\cdots)\big).$
Nested risk measures are useful because they ensure time consistency. That is, the risk evaluation at earlier times is consistent with future evaluations. One of the drawbacks of nested risk measures is that they are typically hard to interpret. We note that typically in the RL literature, the terms ‘dynamic risk measure’ and ‘nested risk measure’ are used interchangeably; however, formally speaking, dynamic risk measures need not be time-consistent, nor have the nested structure.
@@ Appendix B Numerical Experiments
$\text{CVaR}_{\pi_{t}}\doteq\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=t_{0}}^{n}\text{CVaR}[R_{t}\mid S_{t}\sim\mu_{\pi_{t}},A_{t}\sim\pi_{t}].$ (B.1)
@@ B.1 $\tau$-RPBP Task
In the first task, we considered a continual variation of the RPBP task, such that the risk attitude of the agent, which is governed by the CVaR parameter, $\tau$, changes over time from risk-neutral ($\tau=0.9$) to risk-averse ($\tau=0.1$). In particular, we would expect that the agent first learns to stay in the blue world state, but then changes its preference to the red world state as its risk attitude changes from risk-neutral to risk-averse. More formally, this task can be viewed as a continual learning task, $\{\mathcal{M}_{k}\}_{k=1}^{2}$, with a changing reward function, such that $\mathcal{M}_{k}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R}_{k},p\rangle$, where
@@ B.2 $\mathcal{S}$-RPBP Task
In the second task, we considered another variation of the RPBP task. In this variation, the characteristic per-step reward distributions of the states change over time, such that the agent is required to continually adapt and find the state with the better CVaR (given a fixed risk attitude, $\tau$). More formally, this task can be viewed as a continual learning task with a changing state-space, such that a given state is effectively replaced with a state with a different per-step reward distribution. That is, we have a continual learning task, $\{\mathcal{M}_{k}\}$, such that $\mathcal{M}_{k}\doteq\langle\mathcal{S}_{k},\mathcal{A},\mathcal{R},p\rangle$. In particular, for a given $\mathcal{S}_{k}$, the red world state reward distribution is characterized as a Gaussian distribution with mean, $\mu_{\text{red}}$, and standard deviation, $\sigma_{\text{red}}$. Conversely, the blue world state reward distribution is characterized as a mixture of two Gaussian distributions with means, $\mu_{\text{blue-a}}$ and $\mu_{\text{blue-b}}$, standard deviations, $\sigma_{\text{blue-a}}$ and $\sigma_{\text{blue-b}}$, and a mixing coefficient of 0.5.

### v2

@@ 4.3.1 An Ergodic Reinforcement Learning Objective
As per Definition 4.7, time consistency (as defined in Definition A.5) does not hold for ergodic risk measures. However, we note that it can be easily checked that, under ergodicity-like assumptions, we can recover a weaker notion of time consistency for the risk-aware objective (5) as $t\to\infty$ (that is, the time consistency condition (A.1) holds as $t\to\infty$, but not for all $t$, as is required by Definition A.5). Accordingly, this shows that it is still possible to recover some (weaker) notion of time consistency with ergodic risk measures.
@@ 5 Case Study: CVaR as an Ergodic Risk Measure
In this section, we present a case study in which we show how optimizing an ergodic risk measure can induce sensible risk-aware behaviour in a continual learning setting. That is, we show empirically how optimizing an ergodic risk measure can enable the agent to successfully adapt in situations where either its risk attitude changes, or the observations from the environment change.
To this end, we focus on optimizing the well-known conditional value-at-risk (CVaR) risk measure [Rockafellar and Uryasev, 2000]. More formally, consider a random variable $X$ with a cumulative distribution function, $F(x)=\mathbb{P}(X\leq x)$. The (left-tail) value-at-risk (VaR) of $X$ with parameter $\tau\in(0,1)$ represents the $\tau$-quantile of $X$, such that $\text{VaR}_{\tau}(X)=\sup\{x\mid F(x)\leq\tau\}$. When $F(x)$ is continuous at $x=\text{VaR}_{\tau}(X)$, $\text{CVaR}_{\tau}(X)$ can be interpreted as the expected value of $X$ conditioned on $X\leq\text{VaR}_{\tau}(X)$, such that $\text{CVaR}_{\tau}(X)=\mathbb{E}[X\mid X\leq\text{VaR}_{\tau}(X)]$. Importantly, as per the results in Section 4, and given Assumptions 4.9 and 4.10, the CVaR risk measure can be formulated as the following continual learning objective that corresponds to an ergodic risk measure:
$\text{CVaR}_{\pi_{t}}\doteq\lim_{h\rightarrow{}\infty}\frac{1}{h}\sum_{n=n_{0}}^{h}\text{CVaR}_{\tau}[R_{n}\mid S_{n}\sim\mu_{\pi^{\omega(n_{0})}},A_{n}\sim\pi^{\omega(n_{0})}].$ (6)
In this case study, we optimize the CVaR objective (6) via the RED CVaR Q-learning algorithm proposed in Rojas and Lee [2025] (see Remark 4.12) in two continual learning tasks.
@@
In terms of empirical results, Figures 3 and 4 show the resulting agent behaviour as learning progresses in both tasks. In particular, Figure 3 shows that in the $\tau$-RPBP task, the agent correctly learns to stay in the blue world state in the beginning, and then correctly changes its preference to the red world state once its risk attitude changes from risk-neutral to risk-averse. Similarly, Figure 4 shows that in the $\mathcal{S}$-RPBP task, the agent is able to continually adapt and find the state with the better CVaR. The full set of experimental details and results can be found in Appendix B.
Figure 3: Rolling percent of time that the agent stays in the blue world state as learning progresses in the $\tau$-RPBP task. A solid line denotes the mean percent of time spent in the blue world state, and the shaded region denotes a 95% confidence interval over 50 runs. As shown in the figure, the agent correctly learns to stay in the blue world state in the beginning, and then correctly changes its preference to the red world state once its risk attitude changes from risk-neutral to risk-averse.
Figure 4: Rolling reward CVaR as learning progresses in the $\mathcal{S}$-RPBP task. A solid line denotes the mean CVaR, and the shaded region denotes a 95% confidence interval over 10 runs. The blue and red dashed lines denote the reward CVaR of the blue and red world states, respectively. As shown in the figure, the agent is able to continually adapt and find the state with the better CVaR.
@@ 6 Discussion
In this work, we took the first steps towards developing a risk-aware foundation for continual RL. In particular, we formalized the notion of risk-awareness in a continual setting, and used the resulting framework to propose a set of axioms that establish the conditions needed to enable an agent to be risk-aware in a continual setting. We then examined the classical theory of risk measures, and showed that, in its current form, it is incompatible with continual learning. Then, building on this insight, we extended risk measure theory into the continual setting by introducing a new class of ergodic risk measures that are compatible with continual learning. Finally, we provided a CVaR-based case study, along with numerical results, which showed the intuitive appeal of ergodic risk measures in continual settings.
More broadly, the introduction of ergodic risk measures offers a new alternative for risk-aware learning and decision-making, even in non-continual settings (such as in the non-continual average-reward setting). Moreover, in comparison to the static and nested (dynamic) risk measures that are typically used in non-continual RL settings, we note that ergodic risk measures offer several advantages. In particular, ergodic risk measures retain some notion of time consistency, while remaining highly interpretable, thereby capturing the appeal of both static and nested risk measures.
@@ Appendix A Risk Measures
In essence, a dynamic risk measure provides a time-indexed family of risk assessments, which allows risk to be tracked and updated as new information becomes available. In RL-based contexts, dynamic risk measures can be useful for capturing the sequential nature of decision-making (i.e., that actions taken at each time step can potentially influence future outcomes, and hence, future risk evaluations).
@@
$\rho_{t+1}(X)\leq\rho_{t+1}(X^{\prime})\implies\rho_{t}(X)\leq\rho_{t}(X^{\prime}).$ (A.1)
Time-consistent risk measures can be appealing because they ensure that if one future outcome is deemed less risky than another at some time step, $t$, then that same outcome is not deemed more risky than the other at any other time step. One way to think about time consistency in RL-based contexts is to ask the question: can the agent change its mind about how risky something is based on new information? If the answer is yes, then there exists a lack of time consistency.
@@
(Nested Risk Measure; adapted from Ruszczyński [2010]) A nested risk measure is a dynamic risk measure, $\{\rho_{t}\}_{t=0}^{N}$, that is constructed recursively from one-step conditional risk measures. More formally, given conditional risk measures, $\rho_{t}:L^{\infty}(\mathcal{F}_{t+1})\to L^{\infty}(\mathcal{F}_{t})$, the risk at time $n$, $\rho_{n}$, can be evaluated as:
$\rho_{n}(X)\doteq\rho_{n}\big(\rho_{n+1}(\cdots\rho_{N}(X)\cdots)\big).$ (A.2)
Nested risk measures are useful because they ensure time consistency. That is, the risk evaluation at earlier times is consistent with future evaluations. In RL-based contexts, nested risk measures are also useful because they can induce Bellman-like recursions with appealing dynamic programming-like properties. One of the drawbacks of nested risk measures is that they are typically hard to interpret. We note that typically in the RL literature, the terms ‘dynamic risk measure’ and ‘nested risk measure’ are used interchangeably; however, formally speaking, dynamic risk measures need not be time-consistent, nor have the nested structure.
@@ Appendix B Numerical Experiments
$\text{CVaR}_{\pi_{t}}\doteq\lim_{h\rightarrow{}\infty}\frac{1}{h}\sum_{n=n_{0}}^{h}\text{CVaR}_{\tau}[R_{n}\mid S_{n}\sim\mu_{\pi^{\omega(n_{0})}},A_{n}\sim\pi^{\omega(n_{0})}].$ (B.1)
@@ B.1 $\tau$-RPBP Task
In the first task, we considered a continual variation of the RPBP task, such that the risk attitude of the agent, which is governed by the CVaR parameter, $\tau$, changes over time from risk-neutral ($\tau=0.9$) to risk-averse ($\tau=0.1$). In particular, we would expect that the agent first learns to stay in the blue world state, but then changes its preference to the red world state as its risk attitude changes from risk-neutral to risk-averse. More formally, this task can be viewed as a continual learning task, $\{\mathcal{M}^{k}\}_{k=1}^{2}$, with a changing reward function, such that $\mathcal{M}^{k}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R}_{k},p\rangle$, where
@@ B.2 $\mathcal{S}$-RPBP Task
In the second task, we considered another variation of the RPBP task. In this variation, the characteristic per-step reward distributions of the states change over time, such that the agent is required to continually adapt and find the state with the better CVaR (given a fixed risk attitude, $\tau$). More formally, this task can be viewed as a continual learning task with a changing state-space, such that a given state is effectively replaced with a state with a different per-step reward distribution. That is, we have a continual learning task, $\{\mathcal{M}^{k}\}$, such that $\mathcal{M}^{k}\doteq\langle\mathcal{S}_{k},\mathcal{A},\mathcal{R},p\rangle$. In particular, for a given $\mathcal{S}_{k}$, the red world state reward distribution is characterized as a Gaussian distribution with mean, $\mu_{\text{red}}$, and standard deviation, $\sigma_{\text{red}}$. Conversely, the blue world state reward distribution is characterized as a mixture of two Gaussian distributions with means, $\mu_{\text{blue-a}}$ and $\mu_{\text{blue-b}}$, standard deviations, $\sigma_{\text{blue-a}}$ and $\sigma_{\text{blue-b}}$, and a mixing coefficient of 0.5.